There were 1,700 units in beginning inventory that were 40% complete with regard to conversion. During the month, 8,550 units were started and 9,000 were transferred to finished goods. The ending work in process was 60% complete with regard to conversion costs, and materials are added at the beginning of the process. What is the total amount of equivalent units for materials and conversion at the end of the month using the weighted-average method?
Answer #1

Please find below table useful to compute desired results: -

A B C D E F Equivalent units %DM Direct material % CC Conversion costs completed & transferred 9000 1 =B3*C3 1 =B3*E3 4 endin

End results would be as follows: -

в с D E Equivalent units % DM Direct material % CC Conversion costs 3 completed & transferred | 9,000100% 9,000 100% 9,000 4